25. Your bank offers 3-year certificates of deposit with a stated rate of interest of 16.40% p.a., compounded quarterly. Your cousin (who works at Acme Bank and Trust, a competitor) wants to know what stated rate of interest you would require from them in order to switch your business to their bank. Acme's CD's are compounded on a monthly basis. What is the minimum stated interest rate you should ask for from Acme in order to make you indifferent to switching?
Answer #1

Acme Bank: Stated Rate = 16.4 %, Compounding Frequency: Quarterly

Quarterly Rate = 16.4 / 4 = 4.1 %

Equivalent Monthly Rate = (1.041)^(1/3) - 1 = 0.01348 or 1.348%

Competitor has monthly compounding.

Hence, Required Stated Rate = 1.348 x 12 = 16.181 %